1. A  girl riding her bicycle accelerates from 0 m/s to 8 m/s in 5 seconds. What is the girl’s acceleration?

Formula: a  = (v – v0)

  t




  1.  If a red Corvette, with an initial velocity of 14 m/s, accelerates at a rate of 32 m/s2 in  4 seconds, what will its final velocity, v (use Equation 3)be?

Formula:  v = v0 + at 




  1. A cheetah accelerates from rest with a constant acceleration of 4 m/s2.  How far will it have moved after 12s? 

Formula: distance = speed x time





  1. A rocket, moving with a constant velocity, travels 282 m in 21 s.  What is its velocity?

Formula:  speed = distance/time





1
Expert's answer
2022-01-07T09:29:34-0500

a. The acceleration:


"a=\dfrac{v – v_0}{t}=1.6\text{ m/s}^2."

b. The final velocity:


"v=v_0+at=142\text{ m/s}."

c. The distance:


"d=\dfrac {at^2}2=\dfrac {4·12^2}2=256\text{ m}."


d. The velocity is


"v=d/t=282/21=13.4\text{ m/s}."